# Qubit-Efficient Quantum Annealing for Stochastic Unit Commitment

### v1

@@ IV-A Binary Integer Programming Example
QPHR-ADMM obviates the need for continuous slack variables when handling integer inequality constraints, thereby enabling the optimization problem to achieve optimal solutions entirely within the QPU, without recourse to classical computation. The correctness of the algorithm will be validated by the following example.
$\min_{\boldsymbol{x}\in\{0,1\}^{6}}$ $6x_{1}+3x_{2}-5x_{3}-6x_{4}+4x_{5}-7x_{6}$ (19a)
@@
In this case study, the expression of the objective function follows the structure of the master problem discussed in Section III-B, specifically focusing on integer programming without quadratic terms. Furthermore, all constraints are linear inequality constraints with binary variables. Because there are $6$ binary variables, the qubit string can be denoted as $\left|x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}\right\rangle$.
@@
Theoretically, basic QA requires 6 qubits to solve the problem. However, as there is no continuous variable, each inequality constraint will introduce one slack variable that needs to be encoded by $\mathcal{F}$ binary variables. For the proposed QPHR-ADMM, the default parameters is set as $\eta=1.05$ and $\delta=0.01$, and the entire problem is divided into three ADMM blocks, each containing $2$ qubits. Each block will be iteratively updated in sequence, and upon completion of this process, the coefficients will be updated. The QPU solver, D-Wave Advantage_system 4.1, will independently execute the QA evolution process. In addition, we use the Gurobi solver on a classical computer with an Intel Core i7-12700H processor to obtain the correct target solutions as references.
To test the effectiveness of QPHR-ADMM, three cases are considered: unconstrained, single inequality constraint, and multiple inequality constraints optimization, which can mimic the progressive introduction of Benders cuts. The test results are listed in Table II and Fig. 3 presents the influence of parameters on algorithmic convergence under the activation of two constraints, (19b) and (19c).
First, when no inequality constraints are added, the problem is already in QUBO form. Using a quantum annealer, the correct qubit string solution $\left|001101\right\rangle$ can be obtained through a single complete QPHR-ADMM iteration. Second, by incorporating the constraint (19b) into the original QUBO formulation, the previously obtained solution will no longer satisfy this constraint. Introducing slack variables would require additional qubits for discretization or the involvement of a classical computer to complete the computation. However, utilizing QPHR-ADMM allows us to solve the problem using only $2$ qubits, with the quantum result consistent with those produced by classical computation.
@@
In addition, the multiple inequality constraints optimization emulates the master problem after several iterations of BD, as each iteration of BD creates a new inequality constraint. The results presented in Table II demonstrate that QPHR-ADMM can achieve highly precise solutions for these types of problems. Notably, the ADMM algorithm can reduce the constant qubit occupation from $6$ to $2$, and unlike the basic QA, the qubit overhead will not increase as the number of constraints increases. This capability underscores its potential to effectively address the master problem of SUC.
At last, Fig. 3 confirms the influence of parameters on QPHR-ADMM convergence, as outlined in Section III-C. Smaller penalty coefficients and step size factors ensure stable convergence, while increasing these parameters significantly reduces iteration count, accelerating convergence. For instance, with $\sigma^{0}=0.8$ and $\eta=1.14$, only $5$ iterations are needed. However, overly large coefficients may result in non-convergence. For example, when $\sigma^{0}>0.6$ and $\eta>1.12$, convergence deteriorates sharply, with most cases failing to converge. Therefore, parameter selection requires a trade-off to convergence and speed.
Figure 3: The impact of different coefficients on the number of iterations and convergence (The gray area indicates where the results either fail to converge or converge to incorrect values).
TABLE II: The Optimization Results of (19) Based on QPHR-ADMM.
@@
Constraints Target Solutions $\sigma^{0}$ Quantum Results Iter. Qubit QPHR ADMM Qubit Basic QA
@@
(19b) $011101$ 0.3 $\left|011101\right\rangle$ 3 2 $\mathcal{F}$
(19b) & (19c) $011110$ 0.5 $\left|011110\right\rangle$ 19 2 6+$2\mathcal{F}$
(19b), (19c) & (19d) $110101$ 0.5 $\left|110101\right\rangle$ 4 2 6+$3\mathcal{F}$
IV-B Scenario-based SUC Example
This section explores the feasibility of the QPHR-ADMM method in solving SUC. The power system consists of four generators, an equivalent wind generator, and a load. The output of the wind turbine is influenced by wind speed over a 24-hour period, assuming the wind speed follows a Weibull probability distribution. Additionally, the load demand is uncertain and is assumed to follow a Beta distribution[38].
IV-B1 Theoretical Analysis on Qubit Overhead
Considering that the time horizon of the SUC is $24$ hours, the master problem after BD involves $96$ decision variables for the total commitment statuses. Additionally, $J=12$ and $\mathcal{F}=13$ are set to allocate qubits for discretizing the LB and slack variables, with the corresponding precision coefficients specified as $0.004$ and $0.005$, respectively. Therefore, a total of $108+13k$ qubits are required in the basic QA where $k$ is the number of Bender’s iterations, which is a substantial computational burden for current QPUs. Then, since no binary-encoded slack variables are introduced, the number of qubits required for QPHR-ALM can be determined as $108$. In contrast, based on the QPHR-ADMM algorithm, we decompose the master problem into five unit-based blocks. The first four blocks contain the on///off decision variables for each generator over the $24$-hour period, while the last block is used to compute the discretized LB value. As a result, these five blocks require $24$, $24$, $24$, $24$, and $12$ qubits, respectively, for computation. This implies that the QPHR-ADMM algorithm, based on the D-ADMM framework, requires only a QPU with a capacity of $24$ qubits to finish tasks that would otherwise demand $108+13k$ qubits.
Fig. 4 illustrates how the qubit overhead of the three algorithms in this SUC case study varies with the number of BD iterations. The results indicate that as the number of iterations increases, QPHR-ADMM method demonstrates an increasingly significant advantage in qubit efficiency compared to basic QA.
Figure 4: Qubit overhead of different algorithms with increasing iterations under the BD framwork.
IV-B2 Convergence and Sensitivity Analysis
To start, when we attempt to use basic QA to directly handle the large-scale first Benders iteration, the measured results often fail to converge to the correct ground-state solution, leading to erroneous LB values as purple curves showing in Fig. 5. This failure is primarily attributed to chain break issues [33] in current D-Wave QA-QPUs. These issues arise from the excessive qubit requirements, which sometimes result in some longer chains in embedding. This weakens the chain strength of some qubits, ultimately causing broken chains. Thus, adopting QPHR-ADMM not only reduces qubit overhead for each annealing process, but is also crucial to ensure result accuracy even when the qubit number is efficient.
Table III illustrates the convergence of upper and lower bounds in the BD method under different numbers of scenarios. The UB results are obtained by aggregating all the subproblems, which are processed exclusively by classical computers. The LB can be computed using either QPUs or CPUs. Using quantum methods, BD convergence was ensured after $2$ iterations for a sample of $10$ and $100$ scenarios, respectively. Note that there exits slight discrepancy between the LBs computed by Gurobi and the proposed QPHR-ADMM, due to the precision level of the discretized LB being set at $0.004$.
Fig. 5 shows that the trends in the LB values and the corresponding errors obtained using the QPHR-ADMM method and Basic QA during the first iteration of BD under $100$ sampled scenarios. QPHR-ADMM demonstrates satisfactory convergence performance. For example, as shown in Fig. 5, the LB value gradually narrows down to the optimal value, and the stopping criteria error eventually approaches zero during the QPHR-ADMM iterations. It indicates that the master problem will converge to the correct solution. Additionally, This figure also demonstrates the impact of the initial penalty parameters on convergence performance. For instance, a larger $\sigma^{0}$ can reduce iterations, thereby saving computational time. However, due to the uncertain convergence of QPHR-ADMM, setting a smaller $\sigma^{0}$ is necessary to ensure more reliable convergence, as discussed before.
TABLE III: The BD Results for SUC Problems.
$10$ scenarios $100$ scenarios
00 $27.856$ 00 00 $27.644$ 00 00
$1$ $11.501$ $11.056$ $11.056$ $11.290$ $10.844$ $10.844$
$2$ $11.501$ $11.501$ $11.500$ $11.290$ $11.290$ $11.288$
Figure 5: Convergence of the error (top) and lower bounds (bottom) in the first BD iteration under different values of $\sigma^{0}$.
IV-B3 Promise on the Future Quantum Technique
In the context of QC, its primary potential advantage lies in accelerating computations for power system scheduling. Although current quantum computers are incapable of solving real-scale problems, their potential to handle large-scale problems needs serious consideration. Fig. 6 demonstrates the time required by D-Wave QA processors to solve QUBO problems of varying sizes, focusing solely on direct QPU access time while disregarding queue wait times before and after QPU access. For smaller-scale QUBO problems, classical computers exhibit a clear advantage in computational speed over quantum processors. However, as the problem size increases, the computational time for classical methods escalates significantly due to the NP-hard nature of integer programming. In contrast, the time complexity for quantum computation does not vary a lot. Once the utilization of roughly $100$ equivalent qubits, QC begins to surpass classical methods in efficiency. Nevertheless, due to the risk of chain broken during the embedding process on D-Wave QPUs for larger problems, it is advisable to keep the block size within QPHR-ADMM relatively small.
Figure 6: Average time cost for different size of QUBO problems under a single computation.
V Conclusion
This paper explores the feasibility of solving scenario-based SUC using quantum computation. Within the BD framework, we propose a novel QPHR-ADMM algorithm that builds on QPHR-ALM and D-ADMM. This method enables the master problem to be independently accelerated by a quantum annealing processor, while the subproblems are solved in parallel on CPUs. Additionally, the proposed algorithm effectively mitigates quantum computing failures and significantly reduces qubit overhead, maintaining it as a constant rather than scaling linearly with the number of Benders iterations and quantum precision. By extending the versatility of QA to address a wider range of optimization problems, this study establishes a foundational framework for tackling more complex integer programming problems with inequality constraints in the future.

### v2

@@ V Conclusion
QPHR-ADMM obviates the need for continuous slack variables when handling integer inequality constraints, thereby enabling the optimization problem to achieve optimal solutions entirely within the QPU, without recourse to classical computation. The convergence of the algorithm is first validated by the following example.
$\min_{\bm{x}\in\{0,1\}^{6}}$ $6x_{1}+3x_{2}-5x_{3}-6x_{4}+4x_{5}-7x_{6}$ (21a)
@@
In this case study, the expression of the objective function follows the structure of the master problem discussed in Section III-B, specifically focusing on integer programming without quadratic terms. Furthermore, all constraints are linear inequality constraints with binary variables and the qubit string can be denoted as $\left|x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}\right\rangle$. To test the effectiveness of QPHR-ADMM, three variants of (21) are considered: unconstrained, single inequality constraint, and multiple inequality constraints optimization, which can mimic the progressive introduction of Benders cuts when solving the SUC problem.
@@
Theoretically, basic QA requires 6 qubits for each binary variables, and each inequality constraint will introduce one slack variable that needs to be encoded by $\mathcal{F}$ binary variables. Refering to Table II, the total qubits occupation is $6+k\mathcal{F}$ where $k$ is the number of inequality constraints in this case study and $J$ is omitted as there is no continuous variable. For the proposed QPHR-ADMM, the default parameters are set as $\eta=1.05$ and $\delta=0.01$, and the entire problem is divided into three ADMM blocks, each containing $2$ qubits. These blocks are updated sequentially, and upon completion of this process, the coefficients will be updated.
@@
The test results of QPHR-ADMM are listed in Table V. It demonstrates that QPHR-ADMM can achieve highly precise solutions for all constraints settings. Notably, the ADMM algorithm can reduce the constant qubit occupation from $6$ to $2$, and unlike the basic QA, the qubit overhead will not increase as the number of constraints increases. This capability underscores its potential to effectively address the master problem of SUC.
[…]
At last, QAOA can be integrated into the QPHR-ADMM framework as an alternative QUBO solver. However, as shown in Fig. 7, its constraint satisfaction is inferior to QA. This is attributed to the dispersed solution distribution of QAOA, illustrated in Fig. 8, where the probability of the optimal solution is insufficient to suppress suboptimal outcomes. In the BD framework, such inaccuracies may introduce erroneous Benders cuts and impair subsequent iterations. Therefore, QA remains better suited for QPHR-ADMM, as discussed in Section II-B4.
TABLE V: The Optimization Results of (21) Based on QPHR-ADMM.
@@
Constraints Target Solutions $\sigma^{0}$ Quantum Results Iter. Qubit (QPHR ADMM) Qubit (Basic QA)
@@
[…]
Figure 6: Residual Comparison of QPHR-ALM and QPHR-ADMM.
Figure 7: QAOA against QA: measurement accuracy under different constraint settings and layer depths.
Figure 8: QA and QAOA measurement distributions under constraint (21b) using QPHR-ADMM.